OpenStudy (erniellyd28):

the difference between two numbers is 11. six times the larger number is 17 times the smaller number. write a system of equations describing the given conditions. then solve the system by the substitution method and find the two numbers. 1) let x represent the larger and y why represent the smaller number. type the equaton for "the difference between two numbers is 11" ? 2)type the equation for "six times the larger numer is 17 times the smaller number" ? 3) find the numbers the value of the smaller number is? the value of the larger nyumber is ?

OpenStudy (mathstudent55):

"the difference between two numbers is 11" "let x represent the larger and y why represent the smaller number." Can you write the above as an equation?

OpenStudy (mathstudent55):

What operation does "difference" mean?

OpenStudy (erniellyd28):

subtract?

OpenStudy (mathstudent55):

Correct. That means that you have two numbers you are calling x and y. The difference between them is 11. That means, write an equation where you subtract one number from the other number, and it equals 11. Notice that the difference in 11, a positive number. That means you must subtract the smaller number from the larger number.

OpenStudy (erniellyd28):

21-10 is all i can think of

OpenStudy (mathstudent55):

That is an good example using real numbers because you do get a difference of 11. The problem is telling you about two numbers we don't yet know what they are. We use variables (letters) to represent unknown numbers. The problem specifically tells you to use x for the larger number and y for the smaller number. Can you show in an equation using x and y that the difference between x and y is 11?

OpenStudy (erniellyd28):

x-y=11?

OpenStudy (mathstudent55):

Exactly. You have answered part 1). Note: You wrote it correctly since we know that x is the larger number, and y is the smaller number, x - y will give a positive difference. x - y = 11 is correct. If you had written y - x = 11, that would have been incorrect because when you subtract a larger number from a smaller number, the difference is negative.

OpenStudy (mathstudent55):

Ok, so far we have 1) x - y = 11

OpenStudy (mathstudent55):

Now we need to do part 2)

OpenStudy (mathstudent55):

Remember that the problem told us the larger number is x, and the smaller number is y. Part 2) "type the equation for "six times the larger number is 17 times the smaller number" If the larger number is x, what is 6 times the larger number?

OpenStudy (erniellyd28):

11?

OpenStudy (mathstudent55):

No. We don't yet know what the numbers are. The larger number could be 11, 21, 35, 101, or any other number. For now, we are calling the larger number x. How do you write 6 times the larger number using 6 and x?

OpenStudy (erniellyd28):

x+6=17

OpenStudy (mathstudent55):

x + 6 means "6 added to x", or "6 more than x", or "the sum of x and 6" You are told 6 TIMES the larger number. Which operation, addition, subtraction, multiplication, or division, does "times" mean?

OpenStudy (erniellyd28):

multiplication?

OpenStudy (mathstudent55):

Correct. That means that 6 times the larger number is 6 times x. 6 times x is written as \(6 \times x\), or simply \(6x\) .

OpenStudy (mathstudent55):

For 6 times the larger number we get 6x. Now we need to write 17 times the smaller number. The smaller number is y. Times still means multiplication. How do you write 17 times y in a similar way to 6x?

OpenStudy (erniellyd28):

17y or it can also be written as 17 x Y

OpenStudy (mathstudent55):

Correct. 17 times y = 17 * y = 17y

OpenStudy (mathstudent55):

To finish part 2) we need to put together 6x and 17y. The problem tells us that "six times the larger number is 17 times the smaller number" That means 6x is equal to 17y. Use = and write the equation for 6x and 17y.

OpenStudy (erniellyd28):

this is where i get lost

OpenStudy (mathstudent55):

Why? All the hard work so far is done. We know that 6 times the larger number is 6x. We know that 17 times the smaller number is 17y. To write that 6x is 17y means to simply replace "is" with "=". "is" in the problem's sentence means "is the same as" or "is equal to" The problem is stating "6 times the larger number "IS EQUAL TO" 17 times the smaller number.

OpenStudy (erniellyd28):

soit would be 6x=17?

OpenStudy (mathstudent55):

Close, but not just 17. It's 17 times the smaller number, that we already know is 17y.

OpenStudy (erniellyd28):

6x=17y

OpenStudy (mathstudent55):

Exactly. You have just answered part 2)

OpenStudy (mathstudent55):

So far we have: Part 1) x - y = 11 Part 2) 6x = 17y

OpenStudy (erniellyd28):

this is very hard for me

OpenStudy (erniellyd28):

its very challenging

OpenStudy (mathstudent55):

We're doing it one step at a time.

OpenStudy (mathstudent55):

Now we need to do part 3)

OpenStudy (mathstudent55):

In part 3) we take both equations we wrote in parts 1) and 2) and we solve them as a system of equations using the substitution method.

OpenStudy (erniellyd28):

how would we substitute?

OpenStudy (mathstudent55):

To use the substitution method, we must first solve one equation for one variable. Let's use the first equation, x - y = 11 and solve it for x.

OpenStudy (mathstudent55):

Solving for x means to isolate x. We want x alone on the left side. We have x - y = 11 y is being subtracted from x. In order to get rid of a subtraction, we must do the opposite operation to subtraction which is addition.

OpenStudy (mathstudent55):

In order to undo the subtraction of y, we add y. In an equation, we must do the same operation to both sides, so we add y to both sides of this equation. \(x - y \color{red}{+ y} = 11 \color{red}{+ y}\) Above, I am showing in red that we are adding the same to both sides, y.

OpenStudy (mathstudent55):

We get \(x = 11 + y\) The first equation is now solved for x. Now we are ready to do the substitution step.

OpenStudy (mathstudent55):

Now we take the second equation, \(6x = 17y\) and wehre we see x, we \(\Large \color{red}{substitute} \) what x is equal to.

OpenStudy (mathstudent55):

\(\Large 6 \color{red}{x} = 17y\) \(\Large 6\color{red}{(11 + y)} = 17y\) You see that we had x, but now we have 11 + y instead? That was the substitution step of the substitution method.

OpenStudy (mathstudent55):

Now we have a single equation with only one variable (unknown), the letter y. We can solve for y.

OpenStudy (mathstudent55):

We use the distributive property on the left side, to distribute the 6. \(\Large 6(11 + y) = 17y\) \(\Large 6 \times 11 + 6 \times y = 17y\) \(\Large 66 + 6y = 17y\) Now we want all y's together, so since 6y is being added to 66 on the left side, we subtract 6y from both sides. \(\Large 66 + 6y - 6y = 17y - 6y\) \(\Large 66 = 11y\) We can simply switch sides to get: \(\Large 11y = 66\) We now divide both sides by 11 to get: \(\Large \dfrac{11y}{11} = \dfrac{66}{11} \) \(\Large y = 6\) Since y stands for the smaller number, we now know that the smaller number is 6.

OpenStudy (mathstudent55):

Now we use the first equation: \(\Large x - y = 11\) Since we know y is 6, we substitute 6 for y and solve for x. \(\Large x - 6 = 11\) We add 6 to both siodes: \(\Large x - 6 + 6 = 11 + 6\) \(\Large x = 17\) Now we see that the larger number is 17. The answer is: The two numbers are 17 and 6.

OpenStudy (erniellyd28):

so i would write my last 2 answers in number form?

OpenStudy (mathstudent55):

Finally, we can check the answer to make sure we are correct. First, the difference of the numbers must be 11: 17 - 6 = 11, so that works. Second, 6x = 17y: 6(17) = 17(6) This equation must be true because of the commutative property of multiplication. This equation also checks, and we know for sure that our answer is correct.

OpenStudy (mathstudent55):

1) x - y = 11 2) 6x = 17y 3) 17 and 6

OpenStudy (mathstudent55):

Thew answer to part 3) is 17 and 6

OpenStudy (erniellyd28):

so the value of the smaller number is 6? right?

OpenStudy (erniellyd28):

thank you very very much for your help :-)

OpenStudy (mathstudent55):

Yes, the smaller number is 6, and the larger number is 17. You're welcome.